Phosphorus is an eleneat commonly found in the inf
Answer #1

Answer – there is given reaction –

P4(g) <-----> 2 P2(g)

a)At equilibrium , p(P4) = 0.250 atm , p(P2) = 1.40 atm

we know the Kp expression for this reaction

Kp = p(P2)2 / p(P4)

     = (1.40)2 / 0.250

    = 7.84

b) At initial p(P4) = 1.81*10-3 atm , p(P2) = 0.119 atm

we know reaction quotient , Q = p(P2)2 / p(P4)

                                                = (0.119)2 / 1.81*10-3

                                               = 7.82

So the Kp and Qp are the same, means reaction is at equilibrium (E)

We know

Kp = Kc*(RT)∆n

7.84 = Kc * (0.0821*1210 K)1

Kc = 7.84 / 0.0821*1210

       = 0.0789

We know,

∆Go= -RTlnK

       = -8.314 J/mol.K *1210 K * ln 7.84

       = - 20715 J/mol

       = - 20.7 kJ/mol